OpenStudy (anonymous):

Preform the indicated operations and simplify your answer for the expression: 1/(x^2 -7x -8) - x/(x^2 -12x-32)

OpenStudy (callisto):

First, factor the denominators. (x^2 -7x -8) = x^2 -8x +x - 8 = x(x-8) + (x-8) = ...? (x^2 -12x-32) = Eh?! Is there any typo for this one??

OpenStudy (anonymous):

nope!no typos! i was doing that, factoring the denominators, and then that happened 0.0

OpenStudy (karatechopper):

If there is something wrong in this, then maybe the answer is No Solution?

OpenStudy (anonymous):

that wasnt a multiple choice answer 0.0

OpenStudy (callisto):

If no typo...... (Sorry, I still can't believe it.... Are you sure it is -32 not +32?)

OpenStudy (callisto):

If there are no typos... then.. \[\frac{1}{(x^2 -7x -8)} - \frac{x}{(x^2 -12x-32)} \]\[=\frac{1(x^2 -12x-32)}{(x^2 -12x-32)(x^2 -7x -8)} - \frac{x(x^2 -7x -8)}{(x^2 -12x-32)(x^2 -7x -8)} \]\[=\frac{x^2 -12x-32-x(x^2 -7x -8)}{(x^2 -12x-32)(x^2 -7x -8)} \]And then, simplify it...

OpenStudy (anonymous):

dang it, i was afraid i was going to have to do that. oh well, thanks!

OpenStudy (callisto):

Welcome, and good luck!